OpenStudy (anonymous):

Four of Ethan's family members went on a 3 1/3 road trip. If they all shared the drive time equally, how many hours did each family member drive?

OpenStudy (pfenn1):

Do you know how to divide fractions?\[\frac{3\frac13}4=\frac{\frac{10}3}4=\frac{10}3 \times \frac 14=\]